You and your friends are playing in the swimming pool with a 56-cm-diameter beach ball.
How much force would be needed to push the ball completely under water?(answer in N)
Answer #1

force needed to push the ball completely under water

F = ρF* Vdispl* g

ρF= 1000 kg/m^3

d = 56-cm = .56 m

g = 9.8 m/s^2

Vdispl= (4/3)π * (d/2)^3 = (1/6)π *d^3

Vdispl = (4/3)π * (.56/2)^3 = .09195

F = 1000*.09195*9.8

F = 901.11 N